# Representation Learning in a Decomposed Encoder Design for Bio-inspired Hebbian Learning

### v1

@@ Abstract
Modern data-driven machine learning system designs exploit inductive biases on architectural structure, invariance and equivariance requirements, task specific loss functions, and computational optimization tools. Previous works have illustrated that inductive bias in the early layers of the encoder in the form of human specified quasi-invariant filters can serve as a powerful inductive bias to attain better robustness and transparency in learned classifiers. This paper explores this further in the context of representation learning with local plasticity rules i.e. bio-inspired Hebbian learning . We propose a modular framework trained with a bio-inspired variant of contrastive predictive coding (Hinge CLAPP Loss). Our framework is composed of parallel encoders each leveraging a different invariant visual descriptor as an inductive bias. We evaluate the representation learning capacity of our system in a classification scenario on image data of various difficulties (GTSRB, STL10, CODEBRIM) as well as video data (UCF101). Our findings indicate that this form of inductive bias can be beneficial in closing the gap between models with local plasticity rules and backpropagation models as well as learning more robust representations in general.
Keywords Representation Learning  $\cdot$ Hebbian Learning  $\cdot$ Invariances  $\cdot$ Contrastive Learning
@@ 1 Introduction
Computer vision tasks often involve invariance to nuisance variables such as object pose, size, illumination conditions, etc. In the late 80’s and 90’s, involving expert driven model-based design, this has been achieved by stacking and combining quasi-invariant transformations which guarantee that the output remains unchanged for a range of transformations that are irrelevant to the current context Binford and Levitt (1993); Chin and Dyer (1986).
Modern machine learning uses deep neural networks (DNN) to learn implicit transformations from data instead of explicitly defining invariance requirements upfront Bengio (2009); Krizhevsky et al. (2012); LeCun et al. (2004). DNNs have shown to be highly effective in many application scenarios. Nevertheless due to their black box nature, it is difficult to pinpoint precisely which types of invariances were learned and what image features were leveraged by the network during decision making. Recently, symmetries in neural networks as inductive bias gained attention and ways to discover these have been proposed Moskalev et al. (2022); Liu and Tegmark (2022). In sum, the lack of transparency coupled with the brittleness to adversarial perturbations are some of the main challenges of machine learning design Barredo Arrieta et al. (2020); Carvalho et al. (2019).
One way to gain more understanding into the inner-workings of a neural networks-based system is through decomposable design Lipton (2016). The efforts in the 90’s on quasi-invariance may be seen as a form of decomposition that allows for efficient generation of hypotheses for matching Binford and Levitt (1993). Essentially, the appropriate choice of invariant operators for a given application context enables one to disregard nuisance variables that are irrelevant to the task on hand. This aspect of exploiting generative models defined by the modeler to specify choice of invariant operators and characterizing the statistical properties of the operator as a function of noise models has been discussed in the context of performance characterization frameworks for computer vision Thacker et al. (2008). A promising direction, combining benefits of model-based and deep-learning approaches, is to both leverage knowledge about the task in the form of inductive biases and learn the parts which are difficult to model. Baslamisli et al. (2021), for example, define albedo and shading gradient descriptors using physics-based models and refine the initial estimates using a decomposed albedo/shading encoder-decoder architecture with fusion.
Inductive bias for invariant operator selection may come from physics-based generative models, mathematical arguments or from neuroscience. For example, Luan et al. (2018); Pérez et al. (2020) have illustrated that inductive bias in the early layers of the encoder in the form of human specified (quasi-)invariant filters, e.g. Gabor filters, can serve as a powerful tool to attain a more robust performance in learned classifiers, making the neural network design more transparent and explainable. Parallels can be drawn between the design of decomposable vision systems and brain-inspired vision architectures. The visual system of mammals consists of parallel, hierarchical specialized pathways, interpreting different parts of the input signals Ungerleider (1982). Various proposals for brain-inspired architectures draw inspiration from these structures and present decomposable systems that process multiple cues such as color, shape, texture in parallel with specialized pathways von der Malsburg (2014); Hinton (2021); Hawkins et al. (2019).
Learning in the primate brain is thought to occur in an unsupervised manner according to local plasticity rules Hebb (1949). Previous works are seeking to extend neural networks with local brain-inspired learning rule to replace backpropagation Pogodin and Latham (2020); Lillicrap et al. (2016); Oja (1982). The main motivation of this line of research is understanding how learning happens in the brain, which has the potential to alleviate learning in artificial intelligent systems as well as learn more efficiently with less computational resources Wunderlich et al. (2019). However, the performance of neural networks trained with current bio-inspired rules is usually worse than standard backpropagation networks on the same tasks Bartunov et al. (2018).
In this work, we posit that integrating local plasticity rules in a modular decomposable structure can be beneficial in further closing the gap to networks trained with backpropagation. Inspired by previous works that argue that learning in biological systems does not happen on a blank slate Zador (2019), we explore the combination of these two perspectives: inductive bias in the form of visual invariant operators, motivated from physics and mathematical analyses, for disentanglement and biologically inspired local learning. We present a framework composed of multiple encoder networks each one of them is extended with a different invariant visual descriptor (e.g red-green normalization, local binary pattern operator and dual-tree complex wavelet transform). The neural networks are trained under a contrastive loss in a self-supervised fashion. For video data, we design an additional encoder trained in a self-supervised setting to learn motion sensitive representations. We validate the learned representations with a linear classifier on image (GTSRB, STL10, CODEBRIM) and video (UCF101) data. Our findings provide evidence that better classification of learned representations can be achieved by a judicious choice of inductive biases for decomposition.
@@ 2 Related Work
Disentanglement: Disentanglement is a broad concept, familiar nowadays in the first place due to the discussion on how $\ss$-VAEs can disentangle generative factors of variations in the latent space Higgins et al. (2017, 2022) and when they fail to do so Locatello et al. (2019). More generally though, it can be understood as a decomposition, namely imposing the desired structure in the latent space Mathieu et al. (2019). In this sense, modularization of the generative model by using the knowledge about the data generation process as in Niemeyer and Geiger (2021); Nguyen-Phuoc et al. (2019) can also be called disentanglement. Understanding which causal factors generated the data, as well as the knowledge of the task at hand, determine how best to (pre-)process them, e.g. whether illumination or texture is relevant for the given task. We analyze the classification performance of neural networks, consisting of several encoders, each preceded by a certain preprocessing step using visual descriptors such as LBP, color transformations or wavelets. The data is in this way transformed into a suitable feature space to solve the task(s). Several transformations may be important for the same task, but the benefit of decomposition is that the influence of every transformation can be analyzed separately. Generalizations can hence be made regarding which types of data are to be preprocessed in which way. In neural networks, the choice of these quasi-operators is learned implicitly during training. Explicit design of transform operators requires a versatile domain knowledge about the input data, but by building a decomposition pipeline with different preprocessing operators one could also automatize the process of finding out, what is relevant in the data for the given task.
@@
Combination of Invariant Visual Descriptors and Neural Architectures: Various works attempt to create less opaque and more robust pipelines by integrating well-understood operators into data-driven pipelines. For instance, the SIFT detector, which is quasi-invariant to scale, orientation and illumination, was used in combination with a neural architecture Perronnin and Larlus (2015). Based on the Local-Binary Pattern (LBP) operator an adapted convolutional layer with less parameters was proposed that extracts texture features Juefei-Xu et al. (2017); Lin et al. (2020) Lin et al. (2020). Taking strong inspiration from findings on primate vision, evidence has been put forward that fixing the first layers of a convolutional neural network (CNN) to simulate the image processing of the primate primary visual cortex benefits robustness towards input perturbations and adversarial attacks Dapello et al. (2020). Scattering transforms preserve high frequency information and are invariant to geometric transformations and stable to deformations. Comparable performance on classification benchmarks to pure CNN’s has been shown, while using less parameter and providing more theoretical guarantees Oyallon et al. (2017). Li et al. (2020) improved noise-robustness classification accuracy in CNN by replacing down-sampling layers with discrete wavelet transforms.
The former works focus on a single most robust, or most discriminating operator respectively, and train the data-driven part of the model in a supervised manner. Our work investigates the combination of multiple self-supervised neural encoders, each combined with an invariant operator. Explicitly, we use three quasi-invariants, one texture operator (LBP), one color transformation (rg normalization), as well as, a frequency-related operator (wavelets).
Representation Learning/ Contrastive Learning: Learning generalizable features has been attempted by self-supervised learning, which is concerned with the extraction of patterns from data in the absence of labels. Instead of labels, a certain pretext task, for example image inpainting, image colorization or clustering based on pseudo-labels, can be used to shape the representation space Jing and Tian (2019). Mapping the representations of similar instances closer together in the feature space than dissimilar instances can also directly be achieved by contrastive learning Le-Khac et al. (2020). Data augmentations are often used to generate pairs of such similar images, deciding which invariances are learned. Which augmentations work in which context and why is still subject of research though. Recent works have explored contrastive learning in the context of bio-inspired learning Illing et al. (2021); et al. (2022). In our work, we show that we can improve representation learning by integrating well understood visual descriptors to learn more disentangled representation.
@@
Biologically Inspired Learning: Unlike backpropagation updates, Hebbian plasticity rules are local and depend only on the activations of the pre- and post-synaptic neurons and usually a third modulation factor related to reward or other high level signals. There is also a growing body of works in the literature that leverage local plasticity rules to train deep neural networks Bengio (2009); Illing et al. (2021); Krotov and Hopfield (2019); Nøkland and Eidnes (2019); Pogodin and Latham (2020); Xie and Seung (2003); Millidge et al. (2022). Alternatives for backpropagation include systems that decouple the feedforward and feedback network paths to circumvent some of the biologically implausible aspects of backpropagation Amit (2019); Lillicrap et al. (2016); Crafton et al. (2019); Han and Yoo (2019); Moskovitz et al. (2018); Nøkland (2016).
The imposed constraints on gradient propagation usually come with decrease in downstream task performance. Despite recent advances, scaling beyond simple image classification benchmarks remains an issue for bio-inspired systems Bartunov et al. (2018). There are open questions about the merits of using local rules for training deep learning architectures. Some works showed in fact better data efficiency with local Hebbian training Lagani et al. (2021) or more sparse representations Földiak (1990). But generally, the properties of networks trained with local learning rules, especially in the current age of big data-driven systems, are not well understood.
This paper argues that the performance gap of bio-inspired learning rule can be addressed with appropriate network structures and inductive biases. We analyze the representation learning capacity of our framework on classification tasks of various difficulties including in a multi-label setting and show that the framework variant with local plasticity rules benefits the most from this structure which helps in closing the gap to backpropagation models.
@@ 3.1 Decomposition
We explore the decomposition of the input signal using operators with well established properties: wavelet transform, rg normalization and Local Binary Pattern (LBP) (Figure 1). Each extracted representation is sensitive to different features in the image, hence decomposing into different invariant spaces. These operators serve as inductive biases to learn invariant latent representations. We hypothesize that our approach will help leverage quasi-invariant spaces and learn more robust, stable and generalizable representations, particularly in the context of local plasticity rules where the signal cannot be backpropagated across the whole model.
LBP: Local Binary Pattern (LBP) is a visual discriptor that describes the local texture patterns of an image Ojala et al. (2002). LBP assigns a value to each pixel in an image by comparing its intensity with the values of neighboring pixels and then assigning a binary number based on the comparison. LBP is a popular operator in various application domains like face recognition or texture analysis due to its robustness to illumination changes as well as its computational simplicity Maenpaa (2004); Ahonen et al. (2006). When an image is rotated, the indexes of the different neighbouring pixels change which results in different LBP values. Maenpaa (2004) suggests that by circularly rotating the neighbouring pixels until a minimum value for the binary number is obtained, the LBP values stay the same independent of the current rotation of the input image. We use this version of LBP in our experiments.
RG Normalization: To obtain color values without intensity information, it is possible to transform RGB channels of an image to a normalized rg space Gevers and Smeulders (1999). This is helpful to reduce the effect of varying illumination levels onto the task. The normalized rg values can be computed in the following way:
@@
Wavelet Transform: The third decomposition operator that we consider is Dual Tree Complex Wavelet Transform (DTCWT). It is an enhancement of the discrete wavelet transform (DWT) providing better shift invariance Kingsbury (1998). Unlike Fourier transforms which only capture the frequency, discrete wavelet transforms captures both frequency and location in time. DTCWT bears many similarities to the Morlet transform used in the invariant scattering convolution networks Bruna and Mallat (2013). We refer to Valens (1999) for a detailed theory of different wavelet transforms and their respective advantages.
Motion-sensitive Encoder for Video: For video data, we further extend our framework with an additional encoder (MovNet). MovNet is trained with a self-supervised objective for the discrimination between pseudo-egomotion and scene motion in video. Egomotion can be understood as the motion of the observer or the camera, while scene motion is the motion of the objects in a video scene. Further details about the encoder can be found in section 4.7 and Figure 4.
@@ 3.2 Local Learning with Contrastive Predictive Coding
We consider training the encoders in a contrastive predictive coding (CPC) setting before concatenating their extracted representations for downstream tasks. More details about CPC can be found in this paper Oord et al. (2018). The main idea is that the encoder extracts useful representations from a sequence of inputs (e.g sequence of video frames or a sequence of image patches) by learning to predict the future response (an image patch or video frame) in latent space, while at the same time keeping the prediction as far as possible from the responses to negative examples.
A follow-up paper Illing et al. (2021) interprets this future response as a fixation of gaze or as a saccade movement, shifting the center of gaze. The main intuition is that for continuous input streams, the statistics of consecutive frames should have similar latent representations as long as there is a fixation of the eye movement on the same object/scene. In this biologically inspired setting, Contrastive, Local And Predictive Plasticity (CLAPP) Illing et al. (2021) model uses a Hinge loss to train an encoder layer-wise via a Hebbian-like learning rule. The role of Hinge loss is to contrast representations with each other. We use this bio-inspired CLAPP model with local learning for our experiments. In further accordance with Illing et al. (2021), we will also compare the above Hebbian version with a CPC variant using the same Hinge loss but trained end-to-end. This we will refer to as HingeCPC.
In the case of image data, the sequence needed for self-supervision is created by dividing the images patch-wise Oord et al. (2018); Illing et al. (2021). The patches overlap to simulate a sequence of movements from a single image and are stored in a batch in vertical order of the image. Given a sequence of input patches from several images, the encoder network produces a representation $z^{t}$ for each time step t (for every patch). Each of the patches can be interpreted as a context for other patches, determining whether they stem from a continuous scene. The context could in principle encompass several previous time steps, for simplification here $c^{t}=z^{t}$. Specifically, a context representation $c^{t}$ at time t is used to predict the representations of encoded image patches at future time steps $t+\delta$, up to 5 patches into the future. This means that $c^{t}$ is used to predict the representations $z^{t+\delta}$ for $\delta\in\{1..5\}$, where these representations might be encoded patches from the same image (or frames taken from the same video) - positive examples, or randomly sampled patches from other images within the same batch - negative examples. A linear transformation matrix $W^{pred}$ is used for the prediction:
@@
The obtained predictions for future time steps $z_{pred}^{t}$ are either aligned with the actual representations computed by the encoder $z^{t+\delta}$ for positive examples, or contrasted in case of a negative ones, using a Hinge loss function:
@@
$y^{t}$ is the global modulation signal and can be viewed as the label of a binary classification problem characterized by the Hinge loss. If we have a positive example then $y^{t}=1$ and in the case of a negative example $y^{t}=-1$.
Hinge loss can either be applied to the last layer and trained with backpropagation (HingeCPC) or computed for each layer separately. By applying this loss objective layer-wise, learning can proceed based on local information (CLAPP) without the need for top-down feedback as used in backpropagation. CLAPP’S weight update is Hebbian, which only depends on local information available to that layer and a global modulation signal $y^{t}$.
@@ 4.1 Experimental Setting
Datasets: To evaluate our models, we use primarily two standard classification benchmarks (GTSRB and STL10) and one multi-target classification dataset of concrete defects CODEBRIM as well as an action recogntion video dataset UCF101. GTSRB refers to the German Traffic Sign Recognition Benchmark Stallkamp et al. (2011). It contains images of 43 different traffic sign classes, extracted from video recordings under different light and weather conditions. The street sign appearance makes them easy to identify for humans, because of well-defined discriminative features of color, shape, symbols and text. The dataset contains 39,209 training images and 12,630 test images. We use the same training set for both self-supervised encoder training and supervised classifier training.
The STL-10 dataset is an image recognition dataset for unsupervised feature learning Coates et al. (2011), similar to CIFAR10. However, each class has fewer labeled training examples than in CIFAR-10, but a very large set of unlabeled examples for unsupervised learning. The number of unlabeled images is around 100000 whereas the number of labeled images is around 5000. The test set is composed of 800 images for each of the 10 classes. The classes of objects present in the dataset are: airplane, bird, car, cat, deer, dog, horse, monkey, ship, truck.
CODEBRIM is a multi-target classification dataset of defects on concrete surfaces Mundt et al. (2019) with 5 classes of defects: crack, spallation, efflorescence, exposed bars and corrosion stains. It is possible to have multiple defects in the same image. This is a major difference to other benchmarks. CODEBRIM not only allows us to validate our design in a practical real world setting but also investigate if it is possible to infer multiple objects from the latent space extracted by our self-supervised models. The dataset also contains varied and challenging cases due to the overlapping defects, varied appearance of concrete surfaces, presence of surface markings and varied weather conditions. The training set contains 5354 defect bounding boxes with at least one defect and 2506 background bounding boxes of 30 unique bridges, acquired at different scales at a high resolutions. The test set contains 150 defect examples per class.
UCF101 is a standard action recognition dataset to evaluate 3D models and is composed of realistic action videos categorized in 101 action categories Soomro et al. (2012).
Training Details and Hyperparameters: Our encoders are 6-layer-CNNs trained in parallel (see appendix for architecture details). We have two versions of our proposed method: one trained with local learning rule (ours(Local)) and the other one trained with backpropagation (ours(Backprop)).
For comparison, we consider three baseline models: end-to-end supervised classification learning (Supervised-VGG), training with local plasticity rules for contrastive predictive coding (CLAPP) and backpropagation training with Hinge loss formulation for contrastive predictive coding (HingeCPC) Illing et al. (2021). For local training, each layer is trained greedily with the loss objective.
The encoders are implemented in Pytorch and trained on two A100 Nvidia GPUs using Adam with an initial learning rate of $10^{-5}$ and weight decay of $5\cdot 10^{-6}$, mini-batch size 64, where each image is randomly cropped and downsized to 64x64 resolution. We also perform some random flipping of the images with probability $p=0.5$ during training to augment the data. All encoders are trained for 200 epochs. To simulate a sequence of frames for static images, each image is split into patches of size 16x16, similar to Illing et al. (2021); Löwe et al. (2019).
For video data, sequences of 16 frames taken from UCF101 videos are used instead of patches. For training, we use the same hyperparameters as the experiments on image data except a smaller batch size of 16 and 400 instead of 200 epochs. The frames in each video sequence have a resolution of 120x120.
Evaluation Procedure: To quantitatively evaluate the quality of learned representations, we train a linear classifier to recognize the class labels of each dataset while freezing the trained encoder(s). The image patches/video frames are viewed one after the other by the frozen encoders. After that, we perform average pooling over the patch representations to obtain a single representation vector for an input image or video. This vector is then passed to the classifier network. In the case of a multi-encoder setting, the representation vectors of different encoders are concatenated before passing them to the classifier. The linear classifier is composed of two fully connected layers and a dropout layers with $p=0.5$. See appendix for further details.
@@
HingeCPC Illing et al. (2021); Oord et al. (2018) $84.2\pm 0.7$ $69.18\pm 0.07$
CLAPP Illing et al. (2021) $79.61\pm 0.81$ $66.61\pm 0.28$

### v2

@@ Abstract
Modern data-driven machine learning system designs exploit inductive biases in architectural structure, invariance and equivariance requirements, task-specific loss functions, and computational optimization tools. Previous works have illustrated that human-specified quasi-invariant filters can serve as a powerful inductive bias in the early layers of the encoder, enhancing robustness and transparency in learned classifiers. This paper explores this further within the context of representation learning with bio-inspired Hebbian learning rules. We propose a modular framework trained with a bio-inspired variant of contrastive predictive coding, comprising parallel encoders that leverage different invariant visual descriptors as inductive biases. We evaluate the representation learning capacity of our system in classification scenarios using diverse image datasets (GTSRB, STL10, CODEBRIM) and video datasets (UCF101). Our findings indicate that this form of inductive bias significantly improves the robustness of learned representations and narrows the performance gap between models using local Hebbian plasticity rules and those using backpropagation, while also achieving superior performance compared to non-decomposed encoders.
@@ 1 Introduction
Learning in the primate brain is believed to occur in an unsupervised manner, governed by local plasticity rules [16]. Recent research aims to incorporate these brain-inspired learning rules into deep neural networks (DNNs) as an alternative to backpropagation [43, 26, 37]. The main motivation is to understand brain learning mechanisms and develop more computationally efficient systems [48]. However, DNNs trained with current bio-inspired rules often achieve worse performance compared to those trained using standard backpropagation [4].
We propose that integrating local plasticity rules within a modular, decomposable structure can help narrow the performance gap with backpropagation-trained networks. Drawing on the idea that biological learning doesn’t start from a blank slate [50], we explore the synergy between inductive biases, such as visual invariant operators from physics and mathematical analyses, and biologically inspired local learning rules.
DNN implicitly learn transformations from data to solve a certain classification rather than explicitly defining invariances [6, 21, 24]. While DNNs excel in many applications, their black-box nature makes it difficult to identify the learned invariances and features used for decision-making. This lack of transparency, along with vulnerability to adversarial perturbations, presents significant challenges for the safe deployment of systems based on DNNs [3, 9].
@@
To better understand and mitigate these issues, some works have suggested adopting a decomposable design [28]. Computer vision tasks often require invariance to variables such as object pose, size, and illumination. In the late 1980s and 1990s, expert-driven model-based designs achieved this by stacking and combining quasi-invariant transformations, ensuring the output remained stable despite irrelevant changes in context [7, 10].
These efforts highlight the importance of selecting appropriate invariant operators to disregard nuisance variables. This concept, combining model-based and deep learning approaches, leverages task-specific inductive biases while learning parts that are difficult to model [5]. For instance, [29, 41] showed that incorporating human-specified quasi-invariant filters, such as Gabor filters, in the early layers of neural networks can enhance robustness and transparency.
Parallels exist between decomposable vision systems and brain-inspired architectures. The mammalian visual system processes different input signals through parallel, hierarchical pathways [46]. Brain-inspired designs mimic these structures, using specialized pathways to process multiple cues like color, shape, and texture in parallel [32, 17, 15].
We present a framework that incorporates multiple encoder networks, each augmented with distinct invariant visual descriptors such as red-green normalization, local binary pattern operator, and dual-tree complex wavelet transform. These networks are trained using a contrastive loss in a self-supervised manner. To validate the learned representations, we employ a linear classifier on various image datasets (GTSRB, STL10, CODEBRIM) and video datasets (UCF101). For video data, an additional encoder is trained specifically to learn motion-sensitive representations. The video data experiments are in the appendix.
Our findings demonstrate that the careful selection of inductive biases for decomposition significantly enhances the classification of learned representations. This approach is particularly effective with bio-inspired learning rules, helping to bridge the performance gap with backpropagation methods while also achieving better overall performance compared to non-decomposed encoders [^1] Code is accessible here: https://github.com/achrefjaziri/DecomposedEncoders.
@@ 2 Related Work
DNNs with Invariant Operators: Various works aim to create more transparent and robust pipelines by integrating well-understood operators into data-driven systems [40, 25]. For example, the SIFT detector, quasi-invariant to scale, orientation, and illumination, has been combined with neural architectures [42]. Based on the Local Binary Pattern (LBP) operator, an adapted convolutional layer with fewer parameters was proposed to extract texture features [19, 27]. Evidence shows that simulating the primate primary visual cortex (V1) in early CNN layers increases robustness against input perturbations and adversarial attacks [12].
@@
Hebbian Learning: Unlike backpropagation, Hebbian plasticity rules are local, depending only on the activations of pre- and post-synaptic neurons, and often a third factor related to reward or other high-level signals. Recent research leverages these local plasticity rules to train deep neural networks, offering alternatives that decouple feedforward and feedback paths to address the biological implausibility of backpropagation [6, 22, 36, 43, 49, 33, 2, 26, 14, 34].
While these methods often result in decreased performance on downstream tasks, scaling beyond simple image classification remains a challenge for bio-inspired systems [4]. In this work, we show that the performance gap of bio-inspired learning rules can be mitigated with appropriate network structures and inductive biases.
@@ 3.1 Decomposition
We explore the decomposition of the input signal using well-established operators: wavelet transform, rg normalization, and Local Binary Pattern (LBP) (Figure 1). Each extracted representation is sensitive to different image features, creating different invariant spaces. These operators act as inductive biases to learn invariant latent representations. We hypothesize that our approach will leverage quasi-invariant spaces to learn more robust, stable, and generalizable representations, especially in the context of local plasticity rules where the signal cannot be backpropagated across the entire model. An extended version of our pipeline for video data, adding a motion-sensitive encoder, is in the appendix.
LBP: Local Binary Pattern (LBP) is a visual descriptor that captures local texture patterns in an image by comparing each pixel’s intensity with its neighbors and assigning a binary number based on the comparison [38]. LBP is popular in applications like face recognition and texture analysis due to its robustness to illumination changes and computational simplicity [31, 1]. To address rotation sensitivity, [31] suggests circularly rotating the neighboring pixels until the minimum binary value is obtained, ensuring consistent LBP values regardless of image rotation. We employ this rotation-invariant version of LBP in our experiments.
RG Normalization: To obtain color values without intensity information, it is possible to transform RGB channels of an image to a normalized rg space [13]. This is helpful to reduce the effect of varying illumination levels onto the task. The normalized rg values can be computed in the following way:
@@
Wavelet Transform: The third decomposition operator we consider is the Dual Tree Complex Wavelet Transform (DTCWT), an enhancement of the discrete wavelet transform (DWT) that offers better shift invariance [20]. Unlike Fourier transforms, which capture only frequency, DWT captures both frequency and location in time. DTCWT is similar to the Morlet transform used in invariant scattering convolution networks [8]. For detailed theory on different wavelet transforms and their advantages, see [47].
@@ 3.2 Local Learning with Contrastive Predictive Coding
We use Contrastive Predictive Coding (CPC) to train encoders before concatenating their representations for downstream tasks, as detailed in [39]. CPC involves the encoder learning to predict future responses (e.g., image patches or video frames) while distinguishing these from negative examples in latent space.
A follow-up work [18] interprets this future response as eye movement fixations, suggesting that consecutive frames should have similar latent representations when fixating on the same object/scene. In this biologically inspired setting, the Contrastive, Local, And Predictive Plasticity (CLAPP) model uses a Hinge loss to train encoders layer-wise through a Hebbian-like learning rule. We use this CLAPP model for our experiments and compare it with a CPC variant trained end-to-end with Hinge loss, referred to as HingeCPC.
For image data, self-supervision sequences are created by dividing images into overlapping patches, simulating a sequence of movements. The encoder produces a representation $z^{t}$ for each patch. These patches serve as context for others, determining if they stem from a continuous scene. The context $c^{t}$ at time $t$ predicts future representations $z^{t+\delta}$ up to 5 patches ahead. Positive examples come from the same image or video, while negative examples are random patches from other images in the batch. The prediction uses a linear transformation matrix $W^{pred}$:
@@
Predictions $z_{pred}^{t}$ are aligned with actual representations $z^{t+\delta}$ for positive examples or contrasted for negative ones using Hinge loss:
@@
Here, $y^{t}$ is a binary classification label with $y^{t}=1$ for positive and $y^{t}=-1$ for negative examples.
Hinge loss can be applied to the last layer with backpropagation (HingeCPC) or computed layer-wise (CLAPP) for local Hebbian learning.
@@ 4.1 Experimental Setting
Training Details and Hyperparameters: The encoders in our pipeline are 6-layer CNNs trained in parallel with two versions: ours(Local) using local learning rules, and ours(Backprop) using backpropagation.
We compare against three baselines: Supervised-VGG (end-to-end supervised classification), CLAPP (local plasticity rules for CPC), and HingeCPC (backpropagation with Hinge loss for CPC) [18].
Encoders are trained for 200 epochs using Adam with an initial learning rate of $10^{-5}$ and weight decay of $5\cdot 10^{-6}$. We use a mini-batch size of 64, randomly cropping and downsizing images to 64x64, and applying random flipping with probability $p=0.5$. Training lasts for 200 epochs. For image datsets, Static images are split into 16x16 patches to simulate a sequence of frames.
Datasets: We evaluate our models using two standard classification benchmarks (GTSRB [45] and STL10 [11]), one multi-target classification dataset of concrete defects (CODEBRIM [35]), and an action recognition video dataset (UCF101 [44]). Results for the video dataset are in the appendix.
@@
HingeCPC [18, 39] $84.2\pm 0.7$ $69.18\pm 0.07$
CLAPP [18] $79.61\pm 0.81$ $66.61\pm 0.28$